OpenStudy (anonymous):

A laboratory experiment requires 2.25 L of a 1.0 M solution of phosphoric acid (H3PO4), but the only available H3PO4 is a 9.0 M stock solution. How could you prepare the solution needed for the lab experiment?

OpenStudy (btaylor):

Well, use the formula (M1)(V1)=(M2)(V2). You know the required molarity and volume (M1 and V1, respectively), and the initial molarity (M1). You are solving for V1, the volume of the stock solution that you dilute to 2.25 L. V1=(M2)(V2)/(M1), so V1=(1.0)(2.25)/(9.0). By this, V1=0.25 L. So you take 0.25 L of stock solution, add it to 2.00 L of distilled water, and you get your 2.25 L of 1.0 M solution.

OpenStudy (anonymous):

Are you a science teacher or something? haha

OpenStudy (btaylor):

No, I just love and understand chem. I don't know why.

OpenStudy (anonymous):

Dang, lucky! Are you in high school?

OpenStudy (btaylor):

Yes.

OpenStudy (anonymous):

What grade?

OpenStudy (btaylor):

Just finishing my sophomore year.

OpenStudy (anonymous):

Me too! haha I'm in virtual school.